OpenStudy (anonymous):

7. Which conditional and its converse are both true? A. If x = 3, then x2 = 6. B If x = 1, then 2x = 2. C. If x = 2, then x2 = 4. D. If x2 = 4, then x = 2.

OpenStudy (tkhunny):

What do you think? A. If x^2 = 100, then x=10. x = 10 isn't the only solution. B. x^2= 100 if and only if x= 10 or = -10 Not bad. C. If x= -10, then x^2 = 100 That's easy enough to demonstrate. D. if x = 10, then x^2 = 100 That's easy enough to demonstrate. Where does that leave us?

OpenStudy (anonymous):

D.!

OpenStudy (tkhunny):

?? (10)^2 = 100. Noe. That one works.

OpenStudy (anonymous):

B.

OpenStudy (tkhunny):

Why are you guessing? Look at my discussion. See which one works and which doesn't.

OpenStudy (anonymous):

I'm Sorry /.\ All Of Them Works except a because as you said, it is not the only solution

OpenStudy (anonymous):

Answer is A.